OpenStudy (anonymous):

How long will it take a rocket travelling at a rate of 3.9 × 10^6 mph (miles per hour) to travel the 3.8 × 10^8 miles from the earth to the moon? Use the basic physics equation: D = R × t, or Distance = Rate × time.

OpenStudy (anonymous):

hi did you try this one yet :D

OpenStudy (anonymous):

just plug in the given data to the formula D = R × t,

OpenStudy (anonymous):

D = R ( t ) 3.8 × 10^8 miles = 3.9 × 10^6 mph ( t ) find t here

OpenStudy (anonymous):

that is if D = R × t D/R = t

OpenStudy (anonymous):

@clarkexo